# Large Language Models Can Help Mitigate Barren Plateaus

### v1

@@ 2.2 Our Proposed Framework
where $\mathrm{Var}[\partial E^{(t)}]$ denotes the gradient variance in the $t$-th search iteration, and $S^{(t-1)}=\sum_{t_{i}=1}^{t-1}\Delta^{(t_{i})}\cdot I^{(t_{i})}$ denotes the maximum observed gradient variance in the past iterations, where $I^{(t_{i})}$ represents an indicator function $\mathbf{1}\bigl{(}\Delta^{(t_{i})}>\frac{1}{poly(N,L)T}\bigr{)}$ given a condition inside.
@@
Given a certain-size quantum neural network (qnn), the variance of its gradient during training, $\mathrm{Var}[\partial E]$, is bounded by:
@@
From Def. 1 and Lem. 1, during the search of initial model parameters $\bm{\theta}_{0}$ for a certain-size qnn, for $\forall$ $t\in\mathbb{Z}^{+}$, there exist a bound for the expected improvement (EI) s.t.
@@
Lemma 3 (Submartingale Property of EI).
Let $\{\Delta^{(t)}\}_{t\geq 1}$ be an i.i.d. sequence of random variables on a probability space $(\Omega,\mathcal{F},P)$ s.t.
$\begin{cases}P\bigl{(}\Delta^{(t)}>\frac{1}{poly(N,L)T}\bigr{)}=p,\\ P\bigl{(}\Delta^{(t)}\leq\frac{1}{poly(N,L)T}\bigr{)}=1-p,\end{cases}$
for a probability $p\in[0,1]$. We define natural filtration $\mathcal{F}^{(t)}=\sigma\bigl{(}\Delta^{(1)},\Delta^{(2)},\dots,\Delta^{(t)}\bigr{)}$, and the selective accumulation of $\Delta^{(t)}$ for the past $t$ iteration as a stochastic process $\{S^{(t)}\}_{t\geq 1}$ according to Def. 1. Then, $\{S^{(t)}\}_{t\geq 1}$ is a submartingale with respect to the filtration $\{\mathcal{F}^{(t)}\}_{t\geq 1}$.
@@
Figure 3: Analysis of gradient variance trends in the first element of qnns’ model parameters across varying qubit and layer settings for three classic initialization distributions, uniform, normal, and beta. “Classic” denotes that we initialize the model parameters with a classic distribution. “Ours” denotes that we use our framework to search the initial model parameters.
Lemma 4 (Boundedness of Submartingale).
Let $\{S^{(t)}\}_{t\geq 1}$ be a submartingale w.r.t. a $\{\mathcal{F}^{(t)}\}_{t\geq 1}$ s.t. $\sup_{t}\mathbb{E}[|S^{(t)}|]<\infty$. Then, $\{S^{(t)}\}_{t\geq 1}$ is almost surely bounded by a finite constant $B_{S}$ s.t.
$S^{(t)}\leq B_{S},\quad\text{a.s.,}\quad\forall t\in\mathbb{Z}^{+}.$
Dataset $\left|D\right|$ $\left|F\right|$ $\left|C\right|$ Splits
Iris 150 4 3 60:20:20
Wine 178 13 3 80:20:30
Titanic 891 11 2 320:80:179
MNIST 60,000 784 10 320:80:400
3 Experiment
Dataset.
We evaluate our proposed method across four public datasets. Iris is a classic machine-learning benchmark that measures various attributes of three-species iris flowers. Wine is a well-known dataset that includes 13 attributes of chemical composition in wines. Titanic contains historical data about passengers aboard the Titanic and is typically used to predict the survival. MNIST is a widely used small benchmark in computer vision. This benchmark consists of 28$\times$28 gray-scale images of handwritten digits from 0 to 9. We follow the settings of BeInit Kulshrestha and Safro (2022) and conduct experiments in binary classification. Specifically, we sub-sample instances from the first two classes of each dataset to create a new subset. After sub-sampling, we adjust the feature dimensions to ensure they do not exceed the number of available qubits. The statistics of the original datasets, along with the data splits for training, validation, and testing, are presented in Table 1. Importantly, the total number of sub-sampled instances corresponds to the sum of the split datasets. For instance, in the Iris dataset, the total number of sub-sampled instances is 100.
@@ 3 Experiment
Experimental settings.
In the experiment, we analyze the trend of gradient variance by varying the number of qubits ranging from 2 to 20 in increments of 2 (fixed 2 layers) and the number of layers spanning from 4 to 40 in steps of 4 (fixed 2 qubits). To obtain reliable results, we repeat the experiments five times and present them as curves (mean) with their bandwidth (standard deviation). During the search, our framework can identify the optimal model parameters within 50 iterations. In each search iteration, we employ the Adam optimizer with a learning rate of 0.01 and a batch size of 20 to train a qnn with 30 epochs and compute the gradient variance. After training, we compute the expected improvement (EI) and compare it with an assumed lower bound, $\frac{1}{poly(N,L)T}$, in each iteration. We compute the lower bound by $[2^{2N}T]^{-1}$, which is originally designed for uniform distribution. We empirically apply it for all cases as we observe in Fig. 3 that the magnitude of gradient variance is comparable across all datasets.
Evaluation.
We measure the qnn’s training by its gradient variance. A higher gradient variance in training indicates a lower likelihood of being trapped on the barren plateau landscape.
Figure 5: Example of three classic distributions commonly for initialization. In the figure, the red dots represent the initial values of the model parameters.
Searching initial model parameters of qnns via large language models can help alleviate barren plateaus.
We analyze gradient variance trends in the first element of qnns’ model parameters across varying qubit and layer settings for three classic initialization distributions, uniform, normal, and beta distributions, which are presented in Fig. 5 as examples. For each initialization with classic distribution, we compare it (“Classic”) with our proposed methods (“Ours”). As presented in Fig. 3, we observe that in the case of using classic initialization, the gradient variance of qnns will significantly decrease as the number of qubits or layers increases. Compared with it, our method can maintain higher variances, indicating that our framework can mitigate BPs better. In the rest of the experiments, if there is no specific state, we adopt uniform distribution as prior knowledge for posterior estimation.
@@
We further examine whether the content of prompts influences search performance. In the experiments, we tested four prompting scenarios: (i) Including both data description and gradient feedback in prompts (Both desc. and feedback), (ii) Including gradient feedback only (No desc.), (iii) Including data description only (No feedback), (iv) Including neither data description nor gradient feedback (Neither desc. nor feedback). As the results presented in Fig. 4, we observe that suppressing either dataset description or gradient feedback in the prompts leads to a reduction in the gradient variance of qnns. Notably, the reduction is more significant in most cases when gradient feedback is muted compared to the dataset description, suggesting that both factors play a crucial role in mitigating BPs, with gradient feedback contributing significantly more.
llms Acc. Max i/o
GPT 4o 100% 128K/4K
GPT 4o mini 85% 128K/16K
Gemini 1.5 flash 75% 1M/8K
Gemini 1.5 pro 90% 2M/8K
Claude 3.5 sonnet 100% 200K/8K
Table 2: Comparison of initial model parameters’ generation by accuracy (Acc.) using llms, GPT Hurst et al. (2024), Gemini Team et al. (2024), and Claude Anthropic (2024). We measure the generation under different numbers of qubits and layers (20 combinations in total). We also present the maximum number of input/output tokens in the third column.
Comparison of generative performance using llms.
In our proposed framework, the initial model parameters of qnns are generated by llms. In this experiment, we compare the generative performance under varying qnn structures, such as different numbers of qubits or layers. Specifically, we primarily evaluate whether the correct size of model parameters can be generated by testing 20 combinations in accuracy, fixing either 2 layers while varying qubits from 2 to 20 or 2 qubits while varying layers from 4 to 40. As shown in Tab. 2, the results indicate that both GPT-4o and Claude 3.5 Sonnet can achieve 100% accuracy in generating the correct shapes of model parameters. Considering that 4K output tokens are sufficient for our settings, in this study, we mainly use GPT 4o as the backbone llms.
Figure 6: Comparison between two initialization-based strategies, GaInit and BeInit, and our framework, which is initialized with corresponding data distribution for a fair comparison.
@@
We compare our framework with two popular initialization-based strategies, GaInit Zhang et al. (2022) and BeInit Kulshrestha and Safro (2022). Both of them leverage well-designed Normal and Beta distributions to initialize the quantum circuits, respectively. For a fair comparison, we initialize the qnns with the corresponding distribution. We present the results on Iris in Fig. 6 as an example. The results demonstrate that our framework can identify the initial model parameters of qnns that achieve higher gradient variance during training as the model size increases, indicating better mitigation for BPs.

### v2

@@ 4 Our Proposed Framework
where $\mathrm{Var}[\partial E^{(t)}]$ denotes the gradient variance in the $t$-th iteration, and $S^{(t-1)}=\sum_{t_{i}=1}^{t-1}\Delta^{(t_{i})}\cdot I^{(t_{i})}$ denotes the maximum observed gradient variance in the past iterations, where $I^{(t_{i})}$ represents an indicator function $\mathbf{1}\bigl(\Delta^{(t_{i})}\geq\nicefrac{{1}}{{(poly(N,L)K)}}\bigr)$ given a condition inside.
@@
Given a certain-size quantum neural network (qnn), the variance of its gradient at the beginning of training, $\mathrm{Var}[\partial E]$, is bounded by:
@@
From Def. 2 and Lem. 2, in the process of generating initial model parameters $\bm{\theta}_{0}$ for a certain-size qnn, for $\forall$ $t\in\mathbb{Z}^{+}$, there exists a bound for the expected improvement (EI) s.t.
@@
These results indicate that $S^{(t)}$ is $L^{1}$-bounded and integrable for each $t$. Building upon these lemmas, we investigate the submartingale property and rigorously prove in Lem. 4 that $S^{(t)}$ is a submartingale.
Lemma 4 (Informal Statement of Submartingale Property).
Let $\{I^{(t)}\}_{t\geq 1}$ be an i.i.d. sequence of Bernoulli random variables on a probability space $(\Omega,\mathcal{F},P)$ s.t. $I^{(t)}=1$ with a real number $p\in(0,1]$. Then, $\{S^{(t)}\}_{t\geq 1}$ is a submartingale with respect to the filtration $\{\mathcal{F}^{(t)}\}_{t\geq 1}$ which denotes the collections of all possible events up to time $t$.
Note that the random variable is defined in relation to the comparison between the values of $\Delta^{(t)}$ and $\nicefrac{{1}}{{(poly(N,L)K)}}$. From the definition, it is easy to see that $\Delta^{(t)}\geq\nicefrac{{1}}{{(poly(N,L)K)}}$ when $I^{(t)}=1$, and $\Delta^{(t)}<\nicefrac{{1}}{{(poly(N,L)K)}}$ when $I^{(t)}=0$. More precisely, each $\Delta\cdot I$ (omitting the superscripts) defines a joint distribution of one discrete random variable and one continuous random variable. More details about this intuition are provided in the proofs (Appendix A).
Leveraging the convergence of submartingales and the monotonicity of $S^{(t)}$, we establish in Lem. 5 that $S^{(t)}$ has a supremum, which indicates that our proposed framework can eventually generate effective initial model parameters for qnns that can yield non-negligible gradient variance.
Lemma 5 (Boundedness of Submartingale).
Let $\{S^{(t)}\}_{t\geq 1}$ be a submartingale w.r.t. a $\{\mathcal{F}^{(t)}\}_{t\geq 1}$ s.t. $\sup_{t}\mathbb{E}[|S^{(t)}|]<\infty$. Then, $\{S^{(t)}\}_{t\geq 1}$ is almost surely bounded by a finite constant $B_{S}$ s.t. $S^{(t)}\leq B_{S},\quad\text{a.s.,}\quad\forall\ t\in\mathbb{Z}^{+}.$
Building upon Lem. 5, Thm. 4 shows the expected hitting time $\mathbb{E}[T_{b}]$ of a bounded submartingale, ensuring that our framework will converge to a desired solution within a finite number of iterations.
Theorem 4 (Expected Hitting Time of a Bounded Submartingale).
Let $\delta=\nicefrac{{p}}{{(poly(N,L)K)}}>0$ with $p$ defined in Lem. 4. Let $T_{b}$ be the hitting time of a bounded submartingale $\{S^{(t)}\}_{t\geq 1}$, where $S^{(t)}\leq B_{S}$ almost surely, for $\forall\ t\in\mathbb{Z}^{+}$ (by Lem. 5). We define the hitting time as: $T_{b}=\inf\left\{t\in\mathbb{Z}^{+}:S^{(t)}=b\right\},$ for some threshold $b\leq B_{S}$ such that the set is non-empty almost surely. Then the expected hitting time satisfies:
$\mathbb{E}[T_{b}]\leq\frac{b}{\delta}=\frac{b\cdot poly(N,L)\cdot K}{p}.$
With this theorem, it is straightforward to derive the results for two meaningful cases: (i) $b=\nicefrac{{1}}{{poly(N,L)}}$; and (ii) $b=B_{S}$. This is summarized as Cor. 1 in the Appendix A. (i) When the threshold $b=\nicefrac{{1}}{{poly(N,L)}}$, the expected hitting time satisfies $\mathbb{E}[T_{b}]\leq\nicefrac{{K}}{{p}}$, where the probability $p$ is defined in Lem. 4. Note that $K$ determines the desired threshold on increment $\Delta^{(t)}$. This is related to the generative power provided by the input GenAI model $f$, which is manifested in Lem. 4. Concretely, if the model $f$ can help return a wanted $\Delta^{(t)}$ with greater $\nicefrac{{1}}{{(poly(N,L)K)}}$ (i.e. smaller $K$) and better probability (i.e. $p$), then the expected stopping iteration $T$ is getting smaller. (ii) When $b=B_{S}$, i.e., the supremum of the submartingale, $\mathbb{E}[T_{b}]\leq\nicefrac{{B_{S}\cdot poly(N,L)\cdot K}}{{p}}$. This suggests that even when iteratively generating more effective initial parameters for qnns (can yield higher gradient variance), our framework can still converge within a tractable (polynomial) number of iterations if $B_{s}=\mathcal{O}(poly(N,L))$. Both cases demonstrate that our framework is not only theoretically grounded but also robust in generating meaningful initializations for qnns under different optimization goals.
Furthermore, we discuss extreme cases when $p$ is negligible, e.g., $p=\mathcal{O}(2^{-N})$. One such case involves ansatz-induced BPs Holmes et al. (2022), where initialization-based methods fail in this scenario. Another case arises when llms repeatedly generate identical ineffective outputs. The former case could be addressed by modifying the qnn architecture Holmes et al. (2022), while the latter one can be solved by adjusting hyperparameters, like temperature or top-p, to enhance output diversity Achiam et al. (2023).
5 Experiment
Experimental settings.
We evaluate our proposed method across four public datasets, Iris, Wine, Titanic, and MNIST. We present the dataset statistics and settings in the Appendix B. In the experiment, we analyze the trend of gradient variance by varying the number of qubits, ranging from 2 to 20 in increments of 2 (fixed 2 layers), and the number of layers, spanning from 4 to 40 in steps of 4 (fixed 2 qubits). To obtain reliable results, we repeat the experiments five times and present them as curves (mean) with their bandwidth (standard deviation). Overall, our framework can generate effective model parameters within 50 iterations. In each iteration, we employ an Adam optimizer with a learning rate of 0.01 and a batch size of 20 to train a qnn with 30 epochs and compute the gradient variance. After training, we compute the expected improvement (EI) and compare it with an assumed lower bound, $\nicefrac{{1}}{{(poly(N,L)K)}}$, in each iteration, where we set $K=T$ in the experiments. We empirically choose the lower bound by $\nicefrac{{1}}{{(T\cdot N^{6})}}$ (Appendix B) and apply it for all cases, as we observe in Fig. 3 that the magnitudes of gradient variances are comparable across all datasets. For evaluation, we follow McClean et al. (2018) to assess BP by measuring the gradient variance during the early stage of qnns’ training. A higher gradient variance implies a reduced likelihood of encountering BPs.
@@ 5 Experiment
Figure 3: Analysis of gradient variance trends in the first element of qnns’ model parameters across varying qubit and layer settings for three classic initialization distributions, uniform, normal, and beta. “Classic” denotes that we initialize the model parameters with a classic distribution. “Ours” denotes that we use our framework to generate initial model parameters.
Figure 4: Example of three classic distributions commonly used for initialization. In the figure, the red dots represent the initial values of the model parameters.
Generating initial model parameters of qnns using our framework can help mitigate BPs.
We analyze gradient variance trends in the first element of qnns’ model parameters across varying qubit and layer settings for three classic initialization distributions, uniform, normal, and beta distributions, which are presented in Fig. 4 as examples. For each initialization with classic distribution, we compare it (“Classic”) with our proposed methods (“Ours”). As presented in Fig. 3, we observe that in the case of using classic initialization, the gradient variance of qnns will significantly decrease as the number of qubits or layers increases. Compared with it, our method can maintain higher variances, indicating that our framework can mitigate BPs better. In the rest of the experiments, if there is no specific state, we adopt a uniform distribution as prior knowledge for posterior refinement. Besides the above comparison, we further investigate the contribution of llms in the Appendix B.
llms Acc. Max i/o
GPT-4o 100% 128K / 4K
GPT-4o Mini 85% 128K / 16K
Gemini 1.5 Flash 75% 1M / 8K
Gemini 1.5 Pro 90% 2M / 8K
Claude 3.5 Sonnet 100% 200K / 8K
LLaMA 3 70B Instruct 0% 8K / 2K
LLaMA 3 405B Instruct 50% 128K / 2K
Table 1: Comparison of initial model parameters’ generation by accuracy (Acc.) using GPT Hurst et al. (2024), Gemini Team et al. (2024), Claude Anthropic (2024), and LLaMA Grattafiori et al. (2024). Suffix ‘K’ denotes ‘thousand’.
Comparison of generative performance using llms.
In our framework, the initial model parameters of qnns are generated by llms. We compare the generative performance under varying qnn structures, such as different numbers of qubits or layers. Specifically, we primarily evaluate whether the correct size of model parameters can be generated by testing 20 combinations in accuracy, fixing either 2 layers while varying qubits from 2 to 20, or 2 qubits while varying layers from 4 to 40. As shown in Tab. 1, the results indicate that both GPT-4o and Claude 3.5 Sonnet can achieve 100% accuracy in generating the correct shapes of model parameters. Considering 4K output tokens are sufficient for our settings, we mainly use GPT-4o as the backbone llms. We provide additional observation in the Appendix B.
@@
We examine whether the content of prompts influences generative performance. In the experiments, we tested four prompting scenarios: (i) Including both data description and gradient feedback in prompts (Both desc. and feedback), (ii) Including gradient feedback only (No desc.), (iii) Including data description only (No feedback), (iv) Including neither data description nor gradient feedback (Neither desc. nor feedback). As the results presented in Fig. 5, we observe that suppressing either dataset description or gradient feedback in the prompts leads to a reduction in the gradient variance of qnns. Notably, the reduction is more significant in most cases when gradient feedback is muted compared to the dataset description, suggesting that both factors play a crucial role in mitigating BPs, with gradient feedback contributing significantly more.
Figure 6: Comparison between two strategies and our framework, which is initialized with the corresponding data distribution for a fair comparison.
@@
We compare our framework with two representative initialization-based strategies, GaInit Zhang et al. (2022) and BeInit Kulshrestha & Safro (2022). Both of them leverage well-designed Normal and Beta distributions to initialize the qnns, respectively. For a fair comparison, we initialize the qnns with the corresponding distribution. We present the results on Iris in Fig. 6 as an example. The results demonstrate that our framework can generate initial model parameters of qnns that achieve higher gradient variance at the beginning of training as the model size increases, indicating better mitigation for BPs.
Analysis of the expected improvement.
We analyze the patterns on the expected improvement (EI) and the corresponding gradient variance across various qnn structures as iterations progress. Representative experiments conducted on Iris are illustrated in Fig. 11 as an example (Appendix B). Our findings show that the framework can reliably discover meaningful initial parameters regardless of model size. Besides, as the model size grows, more iterations are required to obtain effective initial parameters that enable qnns to maintain higher gradient variance. This is expected, as larger models expand the candidate space, demanding greater computational resources to explore effectively. Both observations verify the Cor. 1 (Appendix A).